OpenStudy (anonymous):

Use transformations to solve the equation. -9b = -45 b = -9 b = -5 b = 5 b = 9

OpenStudy (anonymous):

If you set b = 5, then the left side of the equation will be \[-9*5 = -45\], which is equal to the right side. Then you have solved it. Alternatively:\[-9b = -45 => b = \frac{ -45 }{ -9 } = 5\]

OpenStudy (anonymous):

Remember that \[\frac{ -9b }{ -9 } = \frac{ -45 }{ -9 }\]

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

Glad I could help :)